OpenStudy (anonymous):

how do i set up this equation: the measure of an angle's supplement is 23 degrees more than twice that of its complement. what is the measure of the angle's complement?

jimthompson5910 (jim_thompson5910):

Let x = measure of the unknown angle

jimthompson5910 (jim_thompson5910):

If x is the angle, what is the supplement?

OpenStudy (anonymous):

is it 180?

OpenStudy (anonymous):

complement is 90?

jimthompson5910 (jim_thompson5910):

the angle and its supplement must add to 180

jimthompson5910 (jim_thompson5910):

two complement angles add to 90

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

so if x is the angle, what is the supplement? x + ___ = 180 what goes in the blank?

OpenStudy (anonymous):

46? idk

jimthompson5910 (jim_thompson5910):

let y be the other angle, so x and y are supplements

jimthompson5910 (jim_thompson5910):

this means they add to 180

jimthompson5910 (jim_thompson5910):

x+y = 180

jimthompson5910 (jim_thompson5910):

solve for y: x+y = 180 x+y-x = 180-x y = 180 - x

jimthompson5910 (jim_thompson5910):

So if x is the angle, then 180-x is the supplement (to x)